OpenStudy (anonymous):

Jack usually mows his lawn in 3 hours. Mary can mow the same yard in 5 hours. how much time would it take for them to mow the lawn together? they could mow the lawn in hours if worked together.

OpenStudy (anonymous):

Assuming they split the lawn, it will take Jack 1.5 hours, and Mary 2.5 hours, so together...4 hours. .5 + .5 = 1 1 + 1 + 2 4 hours.

OpenStudy (nobody):

A simple formula for this type of problem is multiply the hours divided by the sum of the hours, i.e. 3*5/(3+5) = 15/8 15/8 hours = 1 7/8 hours combined It does make sense that if Jack could do the job in 3 hours, that if he had help it would take less then 3 hours

OpenStudy (anonymous):

yep nobody is right another way to do this is to assume that the lawn is 150 sq netres in area. i chose 150 because it is divisblr by 3 and 5. Jack aND mary are working together so after i hour Jack has mown 150 / 3= 50 sq meters (1/2 the area) and Mary has done 150 / 5 = 30 that is they have done 80 sq metres in 1 hour. 70 left - so they can do this in 70/80 = 7/8 hour. Total time = 1 7/8 hours.